OpenStudy (anonymous):

Chris Nazarro had total travel expenses of $872. His transportation cost was $136. The hotel cost was $210 and business materials cost $500. If his lunch cost 30% of his dinner, how much was dinner to the nearest cent? Can someone help me with this?

OpenStudy (baru):

how much does chris have left after spending on transport,hotel and business materials?

OpenStudy (anonymous):

26

OpenStudy (baru):

so,26 was lunch+dinner

OpenStudy (baru):

if dinner is 'x' how much is lunch?

OpenStudy (anonymous):

I don't know this is where I got confused haha. Is it .3 ?

OpenStudy (baru):

lunch is 30% of dinner lunch is (30/100 )* x lunch=0.3x

OpenStudy (anonymous):

so does that mean it is 26 / .3

OpenStudy (baru):

we know that lunch + dinner =26 so 0.3x+x=26

OpenStudy (baru):

do you know how to solve that?

OpenStudy (anonymous):

No, not really. :/ Would it be .4x? I don't know

OpenStudy (baru):

no. can you solve if I write it like this? \[\frac{3x}{10} +x=26\]

OpenStudy (baru):

hint: take LCM or common denominator

OpenStudy (anonymous):

1.3x/26 ?

OpenStudy (baru):

you understand that we are trying to find a number for 'x' right

OpenStudy (anonymous):

Yeah, I don't know how to though

OpenStudy (baru):

do you know how to add fractions?

OpenStudy (anonymous):

does x = 20

OpenStudy (baru):

yes it does :)

OpenStudy (baru):

so ur okay with how to find 'x' ?

OpenStudy (anonymous):

Yesssss :) So did I answer my own question or is there more for me to do?

OpenStudy (anonymous):

If lunch + dinner was 26 and x = 20 would the answer be 6 ?

OpenStudy (baru):

the question only asks for price of dinner, x was dinner. if the question asked you to find 'lunch' then that would be 6 like you just showed :)

OpenStudy (anonymous):

Thank you for helping me!! :)

OpenStudy (baru):

no prob :)